The price for a good A has risen from 175 rub. to 210 rub. The demand for a good B has increased from 5400 units to 7100 units. Calculate the cross- price elasticity of demand?


1
Expert's answer
2021-09-25T03:54:41-0400

Cross Elasticity Demand Formula

"E_{BA} = \frac{Percentage \;Change \;in\; Quantity \;of\; B}{Percentage \;Change \;in\; Price \;of\; A} \\

P_A = 210 \\

ΔP = 210-175 = 35 \\

Q_B = 7100 \\

ΔQ = 7100-5400 = 1700 \\

= \frac{\frac{ΔQ_B}{Q_B}}{\frac{ΔP_A}{P_A}} \\

= \frac{\frac{1700}{7100}}{\frac{35}{210}} \\

= \frac{0.2394}{0.1666} \\

= 1.437"